(1/3)log2(x+6) = log83x
Answer by kapilsinghi(68)

(1/3)log2(x+6) = log83x
=> (1/3)log2(x+6) = (log23x)/(log28)
=> (1/3)log2(x+6) = (log23x)/3*(log22)
=> (1/3)log2(x+6) = (1/3)(log23x)
=> log2(x+6) = (log23x)
Taking antilog on both sides
=> (x+6) = 3x
=> 2x = 6
=> x = 3